# Magnon Blockade with Skyrmion Qubit-Magnon Coupling in a Hybrid Quantum System

### v1

@@ Abstract
Magnon blockade (MB) serves as an efficient mechanism for generating a single magnon, which holds significance in quantum information processing. In this work, we propose a scheme to realize magnon blockade effects in a hybrid system consisting of a YIG micromagnet and a skyrmion. Through numerical simulations, we explore the behavior of the second-order correlation function $g^{(2)}(0)$ under various parameter regimes and examine the influence of the ambient thermal noise on the magnon blockade effects. The results indicate that the behavior of the second-order correlation function for the magnon mode critically depends on the skyrmion qubit-magnon coupling strength and the drive detuning. By tuning the drive field, the system effectively suppresses multi-magnon state and yields a high-purity single-magnon state, exhibiting pronounced antibunching and magnon blockade effects. This scheme provides a feasible and versatile new all-magnetic platform for achieving high-purity single-magnon sources.
@@ I Introduction
Hybrid quantum systems leverage the advantages of different quantum subsystems and play a significant role in fields such as quantum information processing, quantum networking, and quantum sensing Wallquist et al. (2009); Kimble (2008a); Xiang et al. (2013); Kurizki et al. (2015). With high spin density, broad frequency tunability, long lifetime, and enhanced coherence times, magnons excited within YIG materials have emerged as promising candidates for quantum information carriers Serga et al. (2010); Chumak et al. (2015); Zheng et al. (2023); Yuan et al. (2022); Rameshti et al. (2022); Kimble (2008b); Huebl et al. (2013); Tabuchi et al. (2014); Li et al. (2019); Wang et al. (2016). The unique properties of magnons have facilitated the development of various hybrid magnonic systems, including cavity optomagnonic systems Chen et al. (2023); Xiong et al. (2023); Zuo et al. (2024), cavity magnonic systems Osada et al. (2016) and qubit-magnon systems Wolski et al. (2020); Trifunovic et al. (2013); Fukami et al. (2021). In these platforms, magnon modes can couple effectively with quantum information carriers such as phonon modes, microwave photons, optical photons, and superconducting qubits Zhang et al. (2016); Shen et al. (2022); Mergenthaler et al. (2017); Xu et al. (2023); Osada et al. (2018); Parvini et al. (2020); Kounalakis et al. (2022); Liu et al. (2019); Xie et al. (2020). A variety of nontrivial quantum phenomena have been observed within these configurations, underscoring the potential of hybrid magnonic systems as powerful platforms for probing and controlling complex quantum interactions Xu et al. (2023); Liu et al. (2019); Xie et al. (2020); Lu et al. (2021); Wang et al. (2023); Xiong (2023); Mukhopadhyay et al. (2022); Liu et al. (2023); Ren et al. (2022); Wang et al. (2019); Zhao et al. (2020a); Hei et al. (2021); Yuan et al. (2020); Qi and Jing (2022); Hisatomi et al. (2016); Zhao et al. (2020b); jun Xu et al. (2021); Wu et al. (2021); Wang et al. (2022); Jin and Jing (2023); Yuan and Duine (2020); Yan et al. (2024); Jin and Jing (2024); Wang et al. (2020); Zhu et al. (2021); Lu et al. (2023); Wu et al. (2024); Couteau et al. (2023); Wang et al. (2024); He et al. (2023).Recently, coupled skyrmion qubit - magnon systems have emerged as a compelling area of research Pan et al. (2024a); Li et al. (2022).
@@
Skyrmions, as magnetic nanostructures with particle-like characteristics, display emerging magnetic effects Reichhardt et al. (2022); Everschor-Sitte et al. (2018); Liensberger et al. (2021); Hirosawa et al. (2022). Their most notable attribute is their topological protection Ochoa and Tserkovnyak (2019); Back et al. (2020), which enables skyrmions to withstand external disturbances. This topological protection allows for state manipulation with ultra-low currents, along with the modulation of electric and magnetic fields Psaroudaki and Panagopoulos (2021); Pan et al. (2024a); Hsu et al. (2016); Birch et al. (2024), facilitating logical transitions within a richly anharmonic operational regime and making them robust information carriers Lonsky and Hoffmann (2020); Reichhardt et al. (2022). Consequently, skyrmions hold significant promise for applications in spintronics and quantum computing Bogdanov and Hubert (1994); Rößler et al. (2006); Mühlbauer et al. (2009); Yu et al. (2010); Jonietz et al. (2010); Schulz et al. (2012); Huang and Chien (2012); Nagaosa and Tokura (2013); Fert et al. (2013, 2017); Bhowal and Spaldin (2022). Experimentally, skyrmions can be generated in centrosymmetric crystals, such as Gd2PdSi3 Kurumaji et al. (2019). In frustrated magnets, skyrmions exhibit internal degrees of freedom related to helicity Leonov and Mostovoy (2015); Lin and Hayami (2016); Kurumaji et al. (2019); Zhang et al. (2017); Yao et al. (2020); Xia et al. (2023). By quantizing collective helicity coordinates, two types of qubits can be constructed Psaroudaki and Panagopoulos (2021), and with coherence times on the order of microseconds, skyrmion qubits are well-suited for quantum state manipulation. Furthermore, nonvolatile readout methodologies provide a robust means for reliable qubit state measurement. In magnon-skyrmion qubit systems, tunable coherent and dissipative coupling between magnon modes in a YIG micromagnet and the quantized helicity degree of freedom of skyrmions has been achieved Pan et al. (2024a). This robust all-magnetic coupling mechanism acts as a mediator to enable long-range coherent interactions with skyrmions, thereby realizing the scalability of skyrmion qubits and providing a promising all-magnetic alternative platform for exploring quantum effects.
@@
In recent years, magnon blockade (MB) has attracted special attention as a purely quantum effect Wang et al. (2024); Xu et al. (2023); Liu et al. (2019); Xie et al. (2020); Zhao et al. (2020b); jun Xu et al. (2021); Wu et al. (2021); Wang et al. (2022); Jin and Jing (2023); Yuan and Duine (2020); Yan et al. (2024); Jin and Jing (2024). The MB effect was first introduced in the context of a hybrid ferromagnet-superconductor quantum system in 2019 Liu et al. (2019). The standard indicator for verifying a stable single-magnon state is $g^{(2)}(0)\rightarrow 0$ Rabl (2011); Liu et al. (2010); Zheng et al. (2019). Analogous to photon blockade Wang et al. (2020); Zhu et al. (2021); Lu et al. (2023); Wu et al. (2024); Couteau et al. (2023); Wu et al. (2024); Lu et al. (2022); Flayac and Savona (2017) and phonon blockade Xie et al. (2017); Xu et al. (2016), magnon blockade enables the preparation of single-magnon states, offering the possibility of quantum control at the single-magnon level.
In this paper, we aim to generate a stable and high-fidelity single-magnon state by optimizing the MB effect, where a YIG micromagnet are coupled to a resonant skyrmion qubit via strong interaction Pan et al. (2024a). The focus of this work is to systematically investigate the steady-state statistical properties of magnons in a skyrmion qubit-magnon coupled system. By numerically solving the quantum master equation for the systems density matrix Liu et al. (2019); Carmichael (1999), we analyze the second-order correlation function $g^{(2)}(0)$, which characterizes the magnon excitation statistics. Our results reveal that $g^{(2)}(0)$ can reach the order of $10^{-6}$, indicating a strong the MB effect. This effect can be optimized within a specific range of system parameters. Furthermore, we demonstrate that the MB effect is highly sensitive to the coupling strength between the skyrmion qubit and magnon, as well as to the detuning of the drive field. We also discuss in detail the impact of thermal noise from the systems environment on the MB effect. Beyond its fundamental scientific significance, the exploration of the MB effect is crucial for the study of magnons at the quantum level, and it opens up avenues for designing single-magnon emitters. Thus, Our findings provide a theoretical foundation for achieving single-magnon quantum control within an all-magnetic platform Ghazaryan et al. (2018).
@@ II Model and HAMILTONIAN
Figure 1: Schematic representation of the hybrid magnon-Skyrmion system.The magnon is probed by an external microwave field $\Omega_{m}$, while the skyrmion is driven by a microwave field $\Omega_{s}$. The magnon mode interacts with the skyrmion mode through a coupling strength $g_{ms}$. The dissipative processes for both the magnon and skyrmion modes are represented by the damping rates $\kappa_{m}$ and $\kappa_{s}$, respectively. Long-lasting spin wave modes are induced by the application of a uniform bias magnetic field $B_{m}$. Here, we concentrate on the Kittel mode, characterized by the coherent precession of all spins in the micromagnet, with spins maintaining identical phase and amplitude.
@@
The physical system considered in our study is a hybrid quantum structure [as illustrated in Fig. 1], which integrates two collective excitation modes: the magnon mode within a ferromagnetic YIG micromagnet and a skyrmion mode positioned beneath it. The strong coupling between them depends on their spatial configuration. Specifically, the vertical distance from the skyrmion to the surface and center of the YIG sphere directly affects the skyrmion qubit-magnon coupling strength, allowing for precise control over the coupling range Pan et al. (2024a). Within the YIG sphere, spin wave modes with long coherence times can be induced through the application of a uniform bias magnetic field $B_{m}$ Pan et al. (2024a); Gonzalez-Ballestero et al. (2020a, b). The magnon represents spin waves in magnetic materials, which are collective excitations of spin ensembles, while the skyrmion is a topological excitation in magnetic materials, typically manifested as rotations or twists in the spin structure. The coupling between the magnon and the skyrmion can be understood as the interaction between the collective spin wave oscillation of the magnon and the localized spin motion within the skyrmion’s topological structure. The free Hamiltonian associated with the magnon mode can be expressed as $\hat{H}_{m}=\omega_{m}\hat{m}^{\dagger}\hat{m}$ (we set $\hbar=1$), where $\hat{m}^{\dagger}$ and $\hat{m}$ represent the creation and annihilation operators of the magnon mode, respectively, and $\omega_{m}=\gamma_{e}B_{m}$ denotes the mode’s resonance frequency, with $\gamma_{e}$ being the gyromagnetic ratio Pan et al. (2024a). By quantizing the collective coordinates, the skyrmion can be treated as a qubit that can be controlled and manipulated. the energy levels of the $\hat{S}_{z}$ skyrmion qubit are anharmonic, the effective Hamiltonian of the $\hat{S}_{z}$ qubit within the subspace $\{|0\rangle,|1\rangle\}$ is given by $\hat{H}_{S_{z}}=\frac{K_{0}}{2}\hat{\sigma}_{z}-\frac{E_{z}}{2}\hat{\sigma}_{x}$ Psaroudaki and Panagopoulos (2021); Pan et al. (2024a); Psaroudaki et al. (2023); Pan et al. (2024b). The interaction Hamiltonian is given by $\hat{H}_{int}=\frac{\tilde{g}_{ms}}{2}(\hat{m}+\hat{m}^{\dagger})\hat{\sigma}_{x}+\frac{g_{ms}}{2}(\hat{m}+\hat{m}^{\dagger})\hat{\sigma}_{z}$an et al. (2024a), where $\tilde{g}_{ms}$ represents the coupling strength along the X and Y axes, and $g_{ms}$ represents the coupling strength along the Z axis.
@@
To achieve and dynamically control magnon blockade, we introduce a microwave drive applied to the skyrmion qubit with a driving frequency $\omega_{s}$. The corresponding Hamiltonian is given by $H_{s}=\Omega_{s}\cos(\omega_{s}t)\hat{\sigma}_{z}$an et al. (2024a). Additionally, to observe the magnon blockade effect, a weak probe field is introduced into the system, with its Hamiltonian described as $H_{d}=\Omega_{d}(\hat{m}^{\dagger}e^{-i\omega_{d}t}+\hat{m}e^{i\omega_{d}t})$ Fujita et al. (2016); Liu et al. (2019), where $\omega_{d}$ and $\Omega_{d}$ are the probe field’s frequency and strength, respectively. The total Hamiltonian of the driven skyrmion qubit-magnon hybrid quantum system, therefore, is given by
$\hat{H}_{\text{total}}=$ $\,w_{m}\hat{m}^{\dagger}\hat{m}+\frac{K_{0}}{2}\hat{\sigma}_{z}-\frac{E_{z}}{2}\hat{\sigma}_{x}$ (1)
@@
$+\Omega_{s}\cos(\omega_{s}t)\hat{\sigma}_{z}+\Omega_{d}(\hat{m}^{\dagger}e^{-i\omega_{d}t}+\hat{m}e^{i\omega_{d}t}).$
When the skyrmion qubit operates near the degeneracy point, its Hamiltonian $\hat{H}_{S_{z}}$ can be expressed in terms of the eigenstates $\lvert\psi_{2}\rangle=\frac{\sqrt{2}}{2}\lvert 1\rangle-\frac{\sqrt{2}}{2}\lvert 0\rangle$ and $\lvert\psi_{1}\rangle=\frac{\sqrt{2}}{2}\lvert 1\rangle+\frac{\sqrt{2}}{2}\lvert 0\rangle$. In the subspace spanned by $\{\lvert\psi_{2}\rangle,\lvert\psi_{1}\rangle\}$, the Pauli operators are defined as $\tilde{\sigma}_{z}=\lvert\psi_{2}\rangle\langle\psi_{2}\rvert-\lvert\psi_{1}\rangle\langle\psi_{1}\rvert$ and $\tilde{\sigma}_{x}=\lvert\psi_{2}\rangle\langle\psi_{1}\rvert+\lvert\psi_{1}\rangle\langle\psi_{2}\rvert$, the coupling term $\tilde{g}_{ms}$ does not induce transitions between energy levels; instead, it causes a minor shift in the energy levels of the skyrmion qubit. This shift is on the order of megahertz ($\tilde{g}_{ms}\approx 50.1$ MHz), which is considerably smaller than the gigahertz-scale resonant frequencies of the hybrid system ($E_{z}\approx\omega_{m}\approx 9.8$ GHz), therefore, this term can be neglected Pan et al. (2024a, b). Using the rotating-wave approximation (RWA) and considering the time-evolution operator in the rotating frame, $U=e^{-i\left(\frac{W_{s}\tilde{\sigma}_{z}t}{2}+w_{d}t\hat{m}^{\dagger}\hat{m}\right)}$, the hybrid quantum system Hamiltonian can be simplified. Thus, the effective Hamiltonian can be written as
$H_{\text{eff}}=$ $\,\Delta_{m}\hat{m}^{\dagger}\hat{m}+\frac{\Delta_{s}}{2}\tilde{\sigma}_{z}+\frac{g_{ms}}{2}(\hat{m}\tilde{\sigma}_{+}+\hat{m}^{\dagger}\tilde{\sigma}_{-})$ (2)
$+\frac{\Omega_{s}}{2}(\tilde{\sigma}_{+}+\tilde{\sigma}_{-})+\Omega_{d}(\hat{m}^{\dagger}+\hat{m}),$
where $\Delta_{m}\equiv\omega_{m}-\omega_{d}$ ($\Delta_{s}\equiv E_{z}-\omega_{s}$) represents the detuning between the magnon mode and the probe field frequency, while $\Delta_{s}$ represents the detuning between the skyrmion qubit frequency and the driving frequency. The approximate conditions are $\omega_{d}\approx\omega_{s}$, $\omega_{d}\gg g_{ms}$, and $\omega_{s}\gg\Omega_{s}$, with $\Delta=\Delta_{s}=\Delta_{m}$.
III Magnon Blockade and the steady-state equal-time second-order correlation function
III.1 the analytical expression for the steady-state equal-time second-order correlation function
In the weak-driving regime, where $\Omega_{d},\Omega_{s}\ll k_{m},k_{s}$, and the effects of quantum jumps on short time scales are neglected, the wave function of the composite system can be approximated within the framework of an analytical model Wu et al. (2021); Liu et al. (2019); Xie et al. (2020); Wang et al. (2022); Kómár et al. (2013); Flayac and Savona (2017); Tang et al. (2019); Jin and Jing (2024, 2023). Under this approximation, the system’s dynamics can be effectively described by the master equation (Eq. (9)), which, in turn, is approximated by a non-Hermitian Hamiltonian. The non-Hermitian Hamiltonian governing the system is given by

### v2

@@ Abstract
Magnon blockade is a fundamental quantum phenomenon for generating single-magnon state, which gradually becomes one of the candidates for quantum information processing. In this paper, we propose a theoretical scheme to generate the magnon blockade in a hybrid system consisting of a YIG micromagnet and a skyrmion. Considering weak probing of the magnon and driving of the skyrmion qubit, the second-order correlation function is analytically derived, and the optimal condition for realizing the magnon blockade is identified. Under the optimal condition, we systematically analyze the behavior of the second-order correlation function $g^{(2)}(0)$ under different parameter regimes. Our analysis shows that with appropriate driving and probing field intensities, the magnon blockade effect can be significantly enhanced, effectively suppressing multi-magnon states and facilitating the generation of high-purity single-magnon states exhibiting pronounced antibunching. Furthermore, we explore the physical mechanisms underlying the magnon blockade, revealing the coexistence and interplay of conventional and unconventional magnon blockade. This scheme provides a versatile all-magnetic platform for generating high purity single-magnon sources.
@@ I Introduction
Hybrid quantum systems combine the advantages of different subsystems, playing a key role in quantum information processing, networking, and sensing Wallquist et al. (2009); Kimble (2008); Xiang et al. (2013); Kurizki et al. (2015). Examples of hybrid systems are magnonic systems Chen et al. (2023); Xiong et al. (2023); Zuo et al. (2024); Osada et al. (2016); Wolski et al. (2020); Trifunovic et al. (2013); Fukami et al. (2021), optomechanical systems Kippenberg and Vahala (2008); Meystre (2013); Aspelmeyer et al. (2014); Bowen and Milburn (2015); Lai et al. (2022), and photonic-qubit systems Zeng et al. (2023), and so on. Magnons excited in YIG materials, with high spin density, frequency tunability, long lifetime, and enhanced coherence, are promising quantum information carriers Serga et al. (2010); Chumak et al. (2015); Zheng et al. (2023); Yuan et al. (2022); Rameshti et al. (2022); Huebl et al. (2013); Tabuchi et al. (2014); Li et al. (2019); Wang et al. (2016); Zhang et al. (2022). The unique properties of magnons have facilitated hybrid magnonic systems like cavity optomagnonic systems Chen et al. (2023); Xiong et al. (2023); Zuo et al. (2024), cavity magnonic systems Osada et al. (2016); Zhang et al. (2023) and qubit-magnon systems Wolski et al. (2020); Trifunovic et al. (2013); Fukami et al. (2021). In these systems, magnon modes effectively couple with quantum information carriers, including phonon modes, microwave photons, optical photons, and superconducting qubits Osada et al. (2018); Zhang et al. (2016); Shen et al. (2022); Mergenthaler et al. (2017); Parvini et al. (2020); Kounalakis et al. (2022); Liu et al. (2019); Xie et al. (2020); Xu et al. (2023). Nontrivial quantum phenomena in these configurations highlight the potential of hybrid magnonic systems as platforms for exploring and controlling complex quantum interactions He et al. (2023a, 2024); Zhang et al. (2024a); Xu et al. (2023); Liu et al. (2019); Xie et al. (2020); Lu et al. (2021); Xiong (2023); Mukhopadhyay et al. (2022); Liu et al. (2023); Ren et al. (2022); Wang et al. (2019); Zhao et al. (2020a); Hei et al. (2021); Yuan et al. (2020); Qi and Jing (2022); Hisatomi et al. (2016); Zhao et al. (2020b); jun Xu et al. (2021); Wu et al. (2021); Wang et al. (2022); Jin and Jing (2023); Yuan and Duine (2020); Yan et al. (2024); Jin and Jing (2024); Wang et al. (2020); Zhu et al. (2021); Lu et al. (2023); Wu et al. (2024); Couteau et al. (2023); Wang et al. (2024); He et al. (2023b).
@@
Skyrmions, as magnetic nanostructures with particle-like characteristics, exhibit novel magnetic effects Reichhardt et al. (2022); Everschor-Sitte et al. (2018); Liensberger et al. (2021); Hirosawa et al. (2022). Their most notable attribute is their topological protection Ochoa and Tserkovnyak (2019); Back et al. (2020), which enables skyrmions to withstand external disturbances. This topological protection enables state manipulation with ultra-low currents and modulation of electric and magnetic fields Psaroudaki and Panagopoulos (2021); Pan et al. (2024a); Hsu et al. (2016); Birch et al. (2024), facilitating logical transitions within a richly anharmonic operational regime and making them robust information carriers Lonsky and Hoffmann (2020); Reichhardt et al. (2022). Consequently, skyrmions hold significant promise for applications in spintronics and quantum computing Bogdanov and Hubert (1994); Rößler et al. (2006); Mühlbauer et al. (2009); Yu et al. (2010); Jonietz et al. (2010); Schulz et al. (2012); Huang and Chien (2012); Nagaosa and Tokura (2013); Fert et al. (2013, 2017); Bhowal and Spaldin (2022). Experimentally, skyrmions can be generated in centrosymmetric crystals, such as Gd2PdSi3 Kurumaji et al. (2019). In frustrated magnets, skyrmions exhibit internal degrees of freedom related to helicity Leonov and Mostovoy (2015); Lin and Hayami (2016); Kurumaji et al. (2019); Zhang et al. (2017); Yao et al. (2020); Xia et al. (2023). By quantizing collective helicity coordinates, two types of qubits can be constructed Psaroudaki and Panagopoulos (2021), with skyrmion qubits achieving microsecond-scale coherence times, making them suitable for quantum state manipulation. Tunable coherent and dissipative coupling has been achieved between YIG magnon modes and the quantized helicity of skyrmions in skyrmion qubit-magnon systems Pan et al. (2024a). This robust all-magnetic coupling mediates long-range coherent interactions between a magnon and skyrmion qubits, thereby realizing the scalability of skyrmion qubits and providing a promising platform for exploring quantum effects.
@@
In particular, magnon blockade (MB) has attracted special attention as a purely quantum effect Wang et al. (2024); Xu et al. (2023); Liu et al. (2019); Xie et al. (2020); Zhao et al. (2020b); jun Xu et al. (2021); Wu et al. (2021); Wang et al. (2022); Jin and Jing (2023); Yuan and Duine (2020); Yan et al. (2024); Jin and Jing (2024). The MB effect arises from two mechanisms: energy-level anharmonicity, which inhibits the excitation of a second magnon after the first magnon is excited to the single-magnon state, termed conventional magnon blockade (CMB) Liu et al. (2019); Xie et al. (2020); Jin and Jing (2023); Yan et al. (2024), and destructive interference between transition pathways to the two-magnon state, known as as unconventional magnon blockade (UMB) Xie et al. (2020). The MB effect has been theoretically studied in hybrid ferromagnet-superconductor systems Liu et al. (2019). Subsequent studies proposed schemes to achieve antibunched light via MB, including quantum-interference-enhanced MB in YIG-superconducting circuits Xie et al. (2020), and tunable MB in ferromagnet-superconductor systems with dual qubits Li et al. (2021), and enhanced MB in optomechanical-magnetic systems using quantum destructive interference among three indirect transition paths Zhang et al. (2024b). The standard indicator for verifying a stable single-magnon state is $g^{(2)}(0)\rightarrow 0$ Rabl (2011); Liu et al. (2010); Zheng et al. (2019). Analogous to photon blockade Wang et al. (2020); Zhu et al. (2021); Lu et al. (2023); Wu et al. (2024); Couteau et al. (2023); Wu et al. (2024); Lu et al. (2022); Flayac and Savona (2017); Huang et al. (2018) and phonon blockade Xie et al. (2017); Xu et al. (2016); Zeng et al. (2021), MB enables the preparation of single-magnon states, offering the possibility of quantum control at the single-magnon level.
In this work, we theoretically study the realization of MB in YIG ferromagnetic insulators. The system comprises a YIG micromagnet strongly coupled to a resonant skyrmion qubit Pan et al. (2024a). Probing and driving fields are applied to the YIG and the skyrmion qubit, with the latter modulated by electric fields. Building on this setup, we analyze the steady-state properties of the coupled system and derive the optimal MB condition, $\Delta/\gamma\approx\pm\frac{1}{2}g_{ms}/\gamma$, which is consistent with numerical simulations Liu et al. (2019); Carmichael (1999). Using the optimal condition, we investigate the effect of the probing-to-driving field strength ratio on the MB effect. At a ratio of approximately 6, the second-order correlation function $g^{(2)}(0)\sim 10^{-7}$, indicating the strongest MB effect. This result underscores the critical role of the probing-to-driving field strength ratio in optimizing the MB effect. Furthermore, we investigate the mechanisms of MB, revealing the coexistence of CMB and UMB. CMB arises from the intrinsic anharmonicity and energy level splitting in the magnon-qubit system, where the non-equidistant energy spectrum inhibits multi-magnon excitations. UMB is induced by quantum interference effects between transition pathways under the influence of both driving and probing fields. Together, these mechanisms ensure the suppression of multi-magnon states and enhance the MB effect. In another scenario, the skyrmion qubit is exclusively modulated by magnetic fields, and the tprobing is restricted to magnons. Here, MB persists with $g^{(2)}(0)\sim 10^{-2}$, and the optimal condition $\Delta/\gamma\approx\pm\frac{1}{2}\tilde{g}_{ms}/\gamma$ remains valid. Therefore, our study provides theoretical support for the generation of a single-magnon source, which is essential for investigating fundamental aspects of quantum magnons and for potential applications in magnon-based quantum simulation and hybrid quantum devices Li et al. (2019); Hou and Liu (2019); Pan et al. (2024a).
@@ II Model and HAMILTONIAN
As shown in Fig.  1, we study a hybrid quantum system, integrating magnon modes in a YIG micromagnet and skyrmion modes beneath it. Strong coupling between the skyrmion qubit and magnon can be tuned by varying their vertical separation and the radius of the YIG sphere, as the coupling strength depends on both factors Pan et al. (2024a). Within the YIG sphere, spin wave modes with long coherence times can be induced through the application of a uniform bias magnetic field $B_{m}$ Pan et al. (2024a); Gonzalez-Ballestero et al. (2020a, b). The magnon represents spin waves, collective excitations of spin ensembles in magnetic materials, while the skyrmion is a topological excitation characterized by rotations or twists in the spin structure. The coupling between the magnon and the skyrmion arises from the interaction between the magnon’s collective spin wave oscillation and the skyrmion’s localized spin motion within its topological structure.
Figure 1: Schematic representation of the hybrid magnon-Skyrmion system.The magnon is probed by an external microwave field $\Omega_{m}$, while the skyrmion is driven by a microwave field $\Omega_{s}$. The magnon mode interacts with the skyrmion mode through a coupling strength $g_{ms}$. The dissipative processes for both the magnon and skyrmion modes are represented by the damping rates $\kappa_{m}$ and $\kappa_{s}$, respectively.
Here, we focus on the Kittel mode, which features the coherent precession of all spins in the micromagnet, with spins maintaining identical phase and amplitude. The free Hamiltonian of the magnon mode is $\hat{H}_{m}=\omega_{m}\hat{m}^{\dagger}\hat{m}$ (with $\hbar=1$), where $\hat{m}^{\dagger}$ and $\hat{m}$ are the creation and annihilation operators of the magnon mode, respectively, and $\omega_{m}=\gamma_{e}B_{m}$ is the resonance frequency, with $\gamma_{e}$ being the gyromagnetic ratio Pan et al. (2024a). By quantizing the collective coordinates, the skyrmion can be treated as a qubit that can be controlled and manipulated. The energy levels of the $\hat{S}_{z}$ skyrmion qubit are anharmonic Psaroudaki and Panagopoulos (2021); Pan et al. (2024a); Psaroudaki et al. (2023); Pan et al. (2024b), the effective Hamiltonian of the $\hat{S}_{z}$ qubit within the subspace $\{|0\rangle,|1\rangle\}$ is $\hat{H}_{S_{z}}=\frac{K_{0}}{2}\hat{\sigma}_{z}-\frac{E_{z}}{2}\hat{\sigma}_{x}$. The parameters $K_{0}$ and $E_{z}$ are externally modulated, with $K_{0}$ controlled by magnetic fields and $E_{z}$ by electric fields, enabling the manipulation of the qubit’s states Psaroudaki and Panagopoulos (2021). The interaction Hamiltonian is $\hat{H}_{\text{int}}=\frac{\tilde{g}_{ms}}{2}(\hat{m}+\hat{m}^{\dagger})\hat{\sigma}_{x}+\frac{g_{ms}}{2}(\hat{m}+\hat{m}^{\dagger})\hat{\sigma}_{z}$, where $\tilde{g}_{ms}$ describes the coupling between the X- and Y-direction magnetic field generated by the YIG sphere and the skyrmion qubit, and $g_{ms}$ describes the coupling between the Z-direction magnetic field from the YIG sphere and the skyrmion qubit. The maximum coupling strength of $\tilde{g}_{ms}$ can reach up to 40 MHz, whereas $g_{ms}$ can achieve up to 200 MHz Pan et al. (2024a).
@@
To achieve and dynamically manipulate MB, we introduce a microwave drive applied to the skyrmion qubit with a driving frequency $\omega_{s}$. The corresponding Hamiltonian is $H_{s}=\Omega_{s}\cos(\omega_{s}t)\hat{\sigma}_{z}$ Pan et al. (2024a). Additionally, to observe the MB effect, a weak probe field is introduced into the system, with Hamiltonian $H_{d}=\Omega_{d}(\hat{m}^{\dagger}e^{-i\omega_{d}t}+\hat{m}e^{i\omega_{d}t})$ Fujita et al. (2016); Liu et al. (2019), where $\omega_{d}$ and $\Omega_{d}$ are the probe field’s frequency and intensity, respectively. The total Hamiltonian is given by
$\hat{H}_{\text{total}}=$ $\,w_{m}\hat{m}^{\dagger}\hat{m}+\frac{K_{0}}{2}\hat{\sigma}_{z}-\frac{E_{z}}{2}\hat{\sigma}_{x}$
@@
$+\Omega_{s}\cos(\omega_{s}t)\hat{\sigma}_{z}+\Omega_{d}(\hat{m}^{\dagger}e^{-i\omega_{d}t}+\hat{m}e^{i\omega_{d}t}).$ (1)
When the skyrmion qubit operates near the degeneracy point, its Hamiltonian $\hat{H}_{S_{z}}$ can be expressed in terms of the eigenstates $\lvert\psi_{2}\rangle=\frac{\sqrt{2}}{2}\lvert 1\rangle-\frac{\sqrt{2}}{2}\lvert 0\rangle$ and $\lvert\psi_{1}\rangle=\frac{\sqrt{2}}{2}\lvert 1\rangle+\frac{\sqrt{2}}{2}\lvert 0\rangle$. In the subspace spanned by $\{\lvert\psi_{2}\rangle,\lvert\psi_{1}\rangle\}$, the Pauli operators are defined as $\tilde{\sigma}_{z}=\lvert\psi_{2}\rangle\langle\psi_{2}\rvert-\lvert\psi_{1}\rangle\langle\psi_{1}\rvert$ and $\tilde{\sigma}_{x}=\lvert\psi_{2}\rangle\langle\psi_{1}\rvert+\lvert\psi_{1}\rangle\langle\psi_{2}\rvert$. The coupling term $\tilde{g}_{ms}$ does not induce energy level transitions but causes a small shift ($\tilde{g}_{ms}\approx 50.1\,\text{MHz}$) in the skyrmion qubit energy levels. Because this shift is much smaller than the gigahertz-scale resonant frequencies ($E_{z}\approx\omega_{m}\approx 9.8\,\text{GHz}$), it can be neglected Pan et al. (2024a, b). Using the rotating-wave approximation (RWA) and considering the time-evolution operator in the rotating frame, $U=e^{-i\left(\frac{\omega_{s}\tilde{\sigma}_{z}t}{2}+w_{d}t\hat{m}^{\dagger}\hat{m}\right)}$, the hybrid quantum system Hamiltonian can be simplified. Thus, the effective Hamiltonian can be written as
$H_{\text{eff}}=$ $\,\Delta_{m}\hat{m}^{\dagger}\hat{m}+\frac{\Delta_{s}}{2}\tilde{\sigma}_{z}+\frac{g_{ms}}{2}(\hat{m}\tilde{\sigma}_{+}+\hat{m}^{\dagger}\tilde{\sigma}_{-})$
$+\frac{\Omega_{s}}{2}(\tilde{\sigma}_{+}+\tilde{\sigma}_{-})+\Omega_{d}(\hat{m}^{\dagger}+\hat{m}).$ (2)
where $\Delta_{m}\equiv\omega_{m}-\omega_{d}$ ($\Delta_{s}\equiv E_{z}-\omega_{s}$) represents the detuning between the magnon mode and the probe field frequency, while $\Delta_{s}$ represents the detuning between the skyrmion qubit frequency and the driving frequency. We choose the parameter conditions as $\omega_{d}\approx\omega_{s}$, $\omega_{d}\gg g_{ms}$, and $\omega_{s}\gg\Omega_{s}$, with $\Delta=\Delta_{s}=\Delta_{m}$, which can be flexibly tuned in the current experimental setup.
III Magnon Blockade and OPTIMAL CONDITIONS
III.1 The solution of the second-order correlation function
In the weak-driving regime, where $\Omega_{d},\Omega_{s}\ll k_{m},k_{s}$, and the effects of quantum jumps on short time scales are neglected, the wave function of the composite system can be approximated within the framework of an analytical model Wu et al. (2021); Liu et al. (2019); Xie et al. (2020); Wang et al. (2022); Kómár et al. (2013); Flayac and Savona (2017); Tang et al. (2019); Jin and Jing (2024, 2023). In particular, the dynamics of the system can be described within an approximate framework using a non-Hermitian Hamiltonian